OpenStudy (anonymous):

Alice is playing a game in which she needs to advance exactly 15 spaces in order to win. The number of spaces that she may advance is determined by rolling a die. What is the probability that she will win in 3 turns?

OpenStudy (mathstudent55):

If you roll a die three times, how many total outcomes are there?

OpenStudy (anonymous):

Um...18?

OpenStudy (mathstudent55):

This is the counting principle. Event 1 has 6 outcomes. Event 2 has 6 outcomes. Event 3 has 6 outcomes. Total number of outcomes = 6 * 6 * 6 = 216.

OpenStudy (anonymous):

Ah...ok.

OpenStudy (anonymous):

But still, I'm a bit confused.

OpenStudy (mathstudent55):

Now we know that by rolling a die three times, there are a total of 216 outcomes. Now we need to find how many outcomes have a sum of 15 for the three rolls of the die.

OpenStudy (anonymous):

That's what I'm confused about; how do you find that out?

OpenStudy (mathstudent55):

Let's figure it out. If one roll gives you a 1, can you get a sum of 15?

OpenStudy (anonymous):

Nope.

OpenStudy (mathstudent55):

What if one roll is 2?

OpenStudy (anonymous):

Still no.

OpenStudy (mathstudent55):

How many ways can you roll a sum of 15? 5 + 5 + 5 = 15 6 + 6 + 3 = 15 6 + 3 + 6 = 15 3 + 6 + 6 = 15 4 + 5 + 6 = 15 4 + 6 + 5 = 15 5 + 4 + 6 = 15 5 + 6 + 4 = 15 6 + 4 + 5 = 15 6 + 5 + 4 = 15

OpenStudy (mathstudent55):

There are 10 ways of getting a sum of 15.

OpenStudy (mathstudent55):

The probability is: P(sum of 15) = 10/216 = 5/108

OpenStudy (anonymous):

Thank you so much for your help on this problem. I know it took up a lot of your time, I'm really sorry. Nonetheless, thank you.

OpenStudy (mathstudent55):

Don't be sorry. You're welcome. I'm glad I was able to help.